OpenStudy (anonymous):

what is product of (z-5)cubed? (I don't know how to write exponents)

OpenStudy (anonymous):

\[(z-5)(z-5)=z^2-10z+25\]

OpenStudy (anonymous):

\[z^2-10z+25(z-5)=z^3-10z^2+25z-5z^2+10z-125\]

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

\[(z^2-10z+25)(z-5)\]

OpenStudy (anonymous):

is what i m eant -125 not -12